OpenStudy (anonymous):

Re : Order from least to greatest 1/3 3/12 9/18

OpenStudy (anonymous):

3/12,1/3,9/18

OpenStudy (anonymous):

3/12=1/4 and 9/18 = 1/2, so you would order it: 3/12, 1/3, 9/18